Given x and y are species whose interaction is governed by x¹=x(-20-x+2y) and y¹=y(-50+x-y) ..... identify the type of interaction 2) determine equilibrium point of model and state possible outcomes of this interaction 3)linearise the system around each equilibrium point and discuss the nature of stability of each equilibrium point....4)sketch phase portrait of the above system

1
Expert's answer
2021-12-27T16:10:36-0500

1)

x' and y' does not depend explicitly on time, it is autonomous system

also, this is nonlinear system


2)

"x(-20-x+2y)=0"

"y(-50+x-y)=0"


equilibrium points:

"(0,0),(0,-50),(-20,0),(120,70)"


Jacobian matrix:

"J=\begin{pmatrix}
 \partial f/ \partial x& \partial f/ \partial y \\
 \partial g/ \partial x & \partial g/ \partial y
\end{pmatrix}=\begin{pmatrix}
 -20-2x+2y & 2x \\
 y& -50+x-2y
\end{pmatrix}"


where

"f(x,y)=x(-20-x+2y)"

"g(x,y)=y(-50+x-y)"


for equilibrium points:

"J(0,0)=\begin{pmatrix}
 -20 & 0 \\
 0 & -50
\end{pmatrix}"


"\begin{pmatrix}
 x' \\
 y' 
\end{pmatrix}=\begin{pmatrix}
 -20 & 0 \\
 0 & -50
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}=\begin{pmatrix}
 -20x \\
 -50y
\end{pmatrix}"


"J(0,-50)=\begin{pmatrix}
 -120 & 0 \\
 -50 & 50
\end{pmatrix}"


"\begin{pmatrix}
 x' \\
 y' 
\end{pmatrix}=\begin{pmatrix}
 -120 & 0 \\
 -50 & 50
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}=\begin{pmatrix}
 -120x \\
 -50x+50y
\end{pmatrix}"


"J(-20,0)=\begin{pmatrix}
 20 & -40 \\
 0 & -70
\end{pmatrix}"


"\begin{pmatrix}
 x' \\
 y' 
\end{pmatrix}=\begin{pmatrix}
 20 & -40 \\
 0 & -70
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}=\begin{pmatrix}
 20x-40y \\
 -70y
\end{pmatrix}"


"J(120,70)=\begin{pmatrix}
 -120 & 240 \\
 70 & -70
\end{pmatrix}"


"\begin{pmatrix}
 x' \\
 y' 
\end{pmatrix}=\begin{pmatrix}
 -120 & 240 \\
 70 & -70
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix}=\begin{pmatrix}
 -120x+240y \\
 70x-70y
\end{pmatrix}"


3)

find eigenvalues:


for "J(0,0)" :

"(-20-r)(-50-r)=0"

"r_1=-20,r_2=-50"

eigenvalues are real, r2 < r2 < 0

so, (0,0) is an asymptotically stable node


for "J(0,-50)" :

"(-120-r)(50-r)=0"

"r_1=-120,r_2=50"

eigenvalues are real, r1 < 0 < r2

so, (0,-50) is a saddle point


for "J(-20,0)" :

"(20-r)(-70-r)=0"

"r_1=20,r_2=-70"

eigenvalues are real, r2 < 0 < r1

so, (0,-50) is a saddle point


for "J(120,70)" :

"(-120-r)(-70-r)-240\cdot70=0"

"r^2+190r-8400=0"

"r=\frac{-190\pm \sqrt{190^2+4\cdot8400}}{2}"

"r_1=37,r_2=-227"

eigenvalues are real, r2 < 0 < r1

so, (120,70) is a saddle point


4)